OpenStudy (juscallmesteve):

Could someone provide me the steps to find this answer? Find f '(a). f(x) = 4x2 − 4x + 3

OpenStudy (juscallmesteve):

My understanding is I should take the derivatives for f(x). (d/dx) (4x^2 - 4x +3) 4*(d/dx)[x^2] - 4*d/dx[x] + d/dx [3] 4*2x-4*1+0 8x-4

OpenStudy (solomonzelman):

Yes, \(\color{#000000 }{ \displaystyle f'(x)=8x-4 }\)

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle f'(a) }\) just means, "the derivative (or, \(\color{#000000 }{ \displaystyle f'(x) }\)) evaluated at \(\color{#000000 }{ \displaystyle x=a }\).

OpenStudy (solomonzelman):

Or, as far as the slope is concerned, \(\color{#000000 }{ \displaystyle f'(a) }\) gives you the instantaneous slope of the function \(\color{#000000 }{ \displaystyle f(x) }\) at \(\color{#000000 }{ \displaystyle x=a }\).

OpenStudy (solomonzelman):

if you have any questions, ask.

OpenStudy (solomonzelman):

I will put up another example, just in case..

OpenStudy (jdoe0001):

Luke, may the power rule be with you but yours is correct Juscallmesteve

OpenStudy (solomonzelman):

\(\color{#0aa000 }{ \displaystyle\bf Problem: }\) \(\color{#000000 }{ \displaystyle f(x)=2x^3-5x^2+5x+3 }\). Find, \(\color{#000000 }{ \displaystyle f'(7) }\). \(\color{#0aa000 }{ \displaystyle\bf Solution: }\) \(\color{#000000 }{ \displaystyle f(x)=2x^3-5x^2+5x+3 }\) \(\tiny \\[0.6em]\) \(\color{#000000 }{ \displaystyle f'(x)=(3-\color{red}{1})2x^{3-\color{red}{1}}-(2-\color{red}{1})5x^{2-\color{red}{1}}+5+0 }\) \(\tiny \\[0.6em]\) \(\color{#000000 }{ \displaystyle f'(x)=(2)2x^{2}-(1)5x^{1}+5 }\) \(\tiny \\[0.6em]\) \(\color{#000000 }{ \displaystyle f'(x)=4x^{2}-5x+5 }\) \(\tiny \\[0.6em]\) \(\color{#000000 }{ \displaystyle f'(\color{blue}{7})=4\color{blue}{(7)}^{2}-5\color{blue}{(7)}+5 }\) \(\tiny \\[0.6em]\) \(\color{#000000 }{ \displaystyle f'(\color{blue}{7})=4\cdot 49-35+5 }\) \(\tiny \\[0.6em]\) \(\color{#000000 }{ \displaystyle f'(\color{blue}{7})=166 }\)